OpenStudy (anonymous):

Yesterday, 20 employees in an office were absent. If these absentees constitutes 25% of the employees how many employees are there?

OpenStudy (anonymous):

Let total no of employees = x 0.25*x = 20 => x = 20/0.25 => x = 80

OpenStudy (anonymous):

u can also do it this way 25% means one-fourth so if one-fourth of something is 20 then other three fourths is 20 + 20 + 20 so total 20 + 20 + 20 + 20 = 80